OpenStudy (anonymous):

I am working on mt homework and having an awful time... would anyone be able to assist me with 19xy(x+y) and explain? Thx

OpenStudy (anonymous):

what you want to know exactly as 19xy(x+y)=19(x^2)y+19x(y^2):)

OpenStudy (anonymous):

Yes, and what method did you use to get this answer - FOIL

OpenStudy (anonymous):

I had 9x(2y) + 19xy(2)

OpenStudy (anonymous):

r u trying to say me fool >( you know that if a(b+c)=a*b+a*c i just used that

OpenStudy (anonymous):

sorry, i'm still not sure what you mean... what do the * symbols mean? Are you multiplying?

OpenStudy (anonymous):

yes multiplying:)

OpenStudy (anonymous):

okay.